# Towards Interpretable Concept Learning over Time Series via Temporal Logic Semantics

### v1

@@ 4.1 Architecture
Our architecture performs interpretable time series classification by integrating symbolic reasoning based on Signal Temporal Logic (STL) with neural representations of trajectories. It operates over a fixed set of human-interpretable STL formulae $\Phi_{C}=\{\varphi_{1},\dots,\varphi_{n}\}$, used as temporal logic concepts. Given an input time series $\bm{x}$, we compute its embedding $\mathcal{H}_{\bm{x}}\in\mathbb{R}^{n}$ by evaluating the robustness $\rho(\varphi_{i},\bm{x})$ for each concept $\varphi_{i}\in\Phi_{C}$. This results in a fully interpretable vector where each dimension reflects the degree to which a known temporal property is satisfied. To assess how informative each concept is for classification, we compute class-specific discriminability scores. The model supports an arbitrary number of classes K, enabling it to scale naturally to multiclass classification tasks. The model supports an arbitrary number of classes $K$, enabling it to scale naturally to multiclass classification tasks. For each class $k$ and formula $\varphi_{i}$, we define a discriminability matrix $\mathcal{G}_{\bm{x}}\in\mathbb{R}^{n\times K}$ with entries
@@ 6 Conclusions
By using the ack environment to insert your (optional) acknowledgements, you can ensure that the text is suppressed whenever you use the doubleblind option. In the final version, acknowledgements may be included on the extra page intended for references.

### v2

@@ 4.1 Architecture
Our architecture performs interpretable time series classification by integrating symbolic reasoning based on Signal Temporal Logic (STL) with neural representations of trajectories. It operates over a fixed set of human-interpretable STL formulae $\Phi_{C}=\{\varphi_{1},\dots,\varphi_{n}\}$, used as temporal logic concepts. Given an input time series $\bm{x}$, we compute its embedding $\mathcal{H}_{\bm{x}}\in\mathbb{R}^{n}$ by evaluating the robustness $\rho(\varphi_{i},\bm{x})$ for each concept $\varphi_{i}\in\Phi_{C}$. This results in a fully interpretable vector where each dimension reflects the degree to which a known temporal property is satisfied. To assess how informative each concept is for classification, we compute class-specific discriminability scores. The model supports an arbitrary number of classes $K$, enabling it to scale naturally to multiclass classification tasks. For each class $k$ and formula $\varphi_{i}$, we define a discriminability matrix $\mathcal{G}_{\bm{x}}\in\mathbb{R}^{n\times K}$ with entries
@@ 6 Conclusions
This study was carried out within the PNRR research activities of the consortium iNEST (Interconnected North-Est Innovation Ecosystem) funded by the European Union Next-GenerationEU (Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS_00000043). This manuscript reflects only the Authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them.